# Legal Fact Prediction: The Missing Piece in Legal Judgment Prediction

### v2

@@ Abstract
Legal judgment prediction (LJP), which enables litigants and their lawyers to forecast judgment outcomes and refine litigation strategies, has emerged as a crucial legal NLP task. Existing studies typically utilize legal facts, i.e., facts that have been established by evidence and determined by the judge, to predict the judgment. However, legal facts are often difficult to obtain in the early stages of litigation, significantly limiting the practical applicability of fact-based LJP. To address this limitation, we propose a novel legal NLP task: legal fact prediction (LFP), which takes the evidence submitted by litigants for trial as input to predict legal facts, thereby empowering fact-based LJP technologies to perform prediction in the absence of ground-truth legal facts. We also propose the first benchmark dataset, LFPBench, for evaluating the LFP task. Our extensive experiments on LFPBench demonstrate the effectiveness of LFP-empowered LJP and highlight promising research directions for LFP. Our code and data are available at https://github.com/HPRCEST/LFPBench.
@@
1Junkai Liu, 3Yujie Tong, 1Hui Huang, 4Bowen Zheng, 5Yiran Hu, 3Peicheng Wu, 2Chuan Xiao, 2Makoto Onizuka, 1Muyun Yangthanks: Correspondence to: Shuyuan Zheng <zheng@ist.osaka-u.ac.jp>, Muyun Yang <yangmuyun@hit.edu.cn>., 2Shuyuan Zheng11footnotemark: 1 1Harbin Institute of Technology, 2Osaka University, 3Zhejiang University, 4Dalian University of Technology, 5Tsinghua University
@@ 1 Introduction
Extensive research efforts have been devoted to achieving accurate LJP. However, most of the existing LJP research is limited to fact-based LJP (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022; Gan et al., 2023)), where the input to the LJP system consists of (ground-truth) legal facts, i.e., facts that are formally established through evidence and determined by the judge. However, the users of LJP, such as litigants and their lawyers, typically confirm their legal facts only at a very late stage of the litigation Medvedeva and Mcbride (2023). Consequently, the application of LJP is largely confined, as the users often seek to predict judgments before litigation or in its early stages to develop and adjust litigation strategies or related plans.
@@
To further facilitate research on LFP and LFP-empowered LJP, this paper introduces the first benchmark dataset, LFPBench, which contains evidence items, legal facts, and judgment outcomes collected from 657 litigation cases in China, covering 10 representative types of civil cases. As such, it can be used to evaluate both the LFP and LJP tasks.
@@
  • First, we propose a novel task, legal fact prediction (LFP), which can empower LJP applications to operate in a wider range of real-world scenarios.
@@ 2.1 Background
In this paper, legal fact refers to the facts of the case determined by the judge during litigation, based on the presentation and cross-examination of evidence by the parties in a legal proceeding Korn (1966). In other words, only facts that can be proven by evidence in a court of law can be acknowledged by the judge as legal facts.
As depicted in Figure 2, in civil law countries such as Germany, France, and China, as well as in common law countries like the UK and the US, court proceedings primarily resolve the following two tasks to reach a judgment:
@@
As evident, legal facts serve as the foundation for the application of law, and before legal fact-finding is complete, it is logically impossible to predict a judgment based on legal facts. In fact, legal facts are usually finalized when the judge reaches a judgment. Therefore, as Medvedeva and Mcbride (2023) has pointed out, utilizing ground-truth legal facts for LJP is impractical. Instead, litigants typically complete evidence collection before litigation or in its early stages, making evidence-based LJP a more reasonable choice. However, existing research on LJP primarily assume the accessibility of legal facts, namely they are all limited to fact-based LJP (e.g., (Luo et al., 2017; Zhong et al., 2018; Chen et al., 2019; Yue et al., 2021; Feng et al., 2022; Wu et al., 2022)), which is mismatched with real-world legal practice.
@@ 2.2 Task Definition
Note that the LFP task is not to summarize the evidence into legal facts. Instead, the available evidence items represent fragmented pieces of legal facts rather than a complete picture. Therefore, the LFP task involves deducing and expanding the evidence into legal facts. Moreover, conflicts or contradictions may exist among evidence items, particularly between the evidence presented by the plaintiff and the defendant. This requires the LFP system to assess the strength and logical coherence of the evidence to resolve the conflicts, making the prediction of legal facts a significant challenge.
Notice the plaintiff’s claims and the evidence list are typically available before the trial, therefore leveraging this information as input for LFP aligns with real-world legal practice.
3 LFPBench
@@
Figure 3: Case type distribution of the LFPBench dataset.
@@ 3.1 Dataset Overview
Existing LJP datasets, such as CAIL2018 Xiao et al. (2018) and ILDC CJPE Malik et al. (2021), do not include the evidence items submitted by litigants, making LFP infeasible. Therefore, we propose the first benchmark dataset for the LFP task, LFPBench. LFPBench consists of data for 657 first-instance cases in China, covering ten types of civil causes of action as shown in Figure 3. Each case includes the plaintiff’s claims, the evidence items submitted by the litigants, ground-truth legal facts, ground-truth judgments for the claims, and more. Therefore, LFPBench can be used for evaluating both LFP and LJP tasks.
The data statistics of LFPBench can be found in Table 1. In LFPBench, defendants present counterevidence in 58.9% of cases, leading to disputes over the determination of legal facts. Consequently, only 38.71% of claims are fully supported, and plaintiffs completely win only 25.27% of the cases. Therefore, predicting the legal facts and judgments of these cases is highly challenging. We also present a data sample of a house lease case in Figure 4, illustrating how conflicting evidence between the litigants complicates the determination of legal facts.
@@ 3.2 Dataset Construction
LFPBench data was extracted from judicial judgments in China. Specifically, we selected ten representative civil litigation causes of action and retrieved 100 written judgments for each type from the China Judgments Online database [^1] https://wenshu.court.gov.cn/. These case types encompass various common disputes over property and personal rights in daily life. To ensure quality, three legal experts reviewed the judgments and excluded those with overly vague descriptions of evidence. Ultimately, 657 cases were retained.
Then, we used regular expressions to extract legal facts and judgment outcomes, as they are typically written in a consistent structure in the documents. Conversely, since the writing format for claims and evidence varies across judgments, we employed GPT-4o to extract this information. Afterward, we conducted a manual review to ensure data integrity.
Finally, our legal experts annotated the judgment outcomes. Specifically, as shown in Figure 4, they aligned each claim with its corresponding judgment and categorized the outcome into three labels based on the level of support: fully supported, partially supported, and rejected. These labels enable us to evaluate the LFP task using classification-based assessment methods. More details are provided in Appendix A.
@@
(a) ChatLaw embeddings.
(b) OpenAI embeddings.
Figure 5: The correlation between the similarity between predicted and ground-truth legal facts and the accuracy of judgment prediction. The embeddings are derived from two different models to enhance credibility.
@@ 4.1 Setup
  • RQ1 (Model & LJP Approach Comparison): How do SOTA models perform in LFP? How do different LJP approaches (evidence-based LJP, fact-based LJP, and LFP-empowered LJP) perform?
  • RQ2 (Challenge & Bias Analysis): What are the challenges of the LFP task? What biases do existing models exhibit when handling the LFP task?
@@
We employ 6 LLMs as the LFP and LJP systems, including the closed-source, general-purpose LLMs GPT-4o-0513 Hurst et al. (2024) and Claude-3.5-Sonnet-20241022222https://docs.anthropic.com/en/docs/about-claude/models, the open-source, general-purpose LLMs Qwen2.5-14B-Instruct Yang et al. (2024) and Llama3.1-Chinese-8B UnicomAI (2024), as well as the open-source legal LLMs LawJustice-Llama3.1-8B BAAI (2024b) and Law-Llama3.1-8B basuo (2024). Additionally, we evaluated other legal LLMs including DISC-LawLLM Yue et al. (2024), Lawyer-Llama-13B-V2 Huang et al. (2023) and AIE-51-8-Law-Model lingminai (2025). However, these models failed to perform the LFP task due to poor instruction-following capabilities, as detailed in Appendix B.2.
@@
We compare the following LJP approaches that differ in their input, with their prompts detailed in Appendix B.3.
@@
Table 5: The proportions of cases in which each model predicts a complete win, a partial win, or a loss for the plaintiff using LFP-empowered LJP. Evidence items are ordered based on their sources, ensuring that all evidence from the plaintiff (or defendant) appears first.
@@
(a) Bias in complete-win rate.
(b) Bias in partial-win rate.
(c) Bias in loss rate.
@@
4.2 Model & LJP Approach Comparison
@@
As shown in Table 2, the closed-source models GPT-4o and Claude3.5 consistently achieve the best performance across different LJP approaches, while the open-source legal LLMs Law-Llama3.1-8B and LawJustice-Llama3.1-8B perform the worst, with accuracy close to random guessing. This discrepancy cannot be attributed solely to the relatively small size of the legal LLMs, as Llama3.1-Chinese-8B, which has the same model size, performs significantly better. One possible explanation is that these legal LLMs are typically fine-tuned on short-text legal QA datasets BAAI (2024a); Yue et al. (2024); Huang et al. (2023), making them less capable of handling complex tasks such as LFP and LJP, which require summarization, reasoning, and deduction over long texts. Future research could explore incorporating general-domain instruction data into fine-tuning to mitigate the catastrophic forgetting of fundamental capabilities. It is also promising to develop more complex and higher-quality reasoning datasets in the legal domain. This would enable legal models to move beyond simple QA tasks and instead leverage their enhanced legal knowledge to analyze problems step by step, ultimately evolving into domain experts that outperform general models.
@@
As shown in Table 2, for the general-purpose LLMs, predicting legal judgments directly from evidence results in a $7.42\%$ decrease in accuracy compared to predictions based on ground-truth legal facts. This indicates that while fact-based LJP research has made considerable progress, its effectiveness heavily relies on the accessibility of legal facts. On the other hand, although the more practice-aligned evidence-based LJP underperforms [^3] Note that for the legal LLMs, evidence-based LJP outperforms the other approaches. However, given their poor performance close to random guessing, this difference is likely due to randomness rather than a meaningful advantage., the incorporation of LFP reduces the performance gap by 38.50%. Therefore, predicting legal facts from evidence first and then making legal judgments based on the predicted facts can significantly improve the accuracy of evidence-based LJP. Compared to fact-based LJP, LFP-empowered LJP accommodates a broader range of LJP scenarios in legal practice, striking a favorable balance between accuracy and applicability.
@@
We convert both the predicted legal facts and the ground truth into embeddings based on ChatLaw-Text2Vec Cui et al. (2023) and OpenAI text-embedding-3-large OpenAI (2024), and then compute their cosine similarity. Figure 5 shows a positive correlation between the similarity of legal facts and the accuracy of legal judgments. This confirms that high-quality predicted legal facts serve as a crucial foundation for precise LJP.
4.3 Challenge & Bias Analysis

### v3

@@ Abstract
Legal judgment prediction (LJP), which enables litigants and their lawyers to forecast judgment outcomes and refine litigation strategies, has emerged as a crucial legal NLP task. Existing studies typically utilize legal facts, i.e., facts that have been established by evidence and determined by the judge, to predict the judgment. However, legal facts are often difficult to obtain in the early stages of litigation, significantly limiting the practical applicability of fact-based LJP. To address this limitation, we propose a novel legal NLP task: legal fact prediction (LFP), which takes the evidence submitted by litigants for trial as input to predict legal facts, thereby empowering fact-based LJP technologies to make predictions in the absence of ground-truth legal facts. We also propose the first benchmark dataset, LFPBench, for evaluating the LFP task. Our extensive experiments on LFPBench demonstrate the effectiveness of LFP-empowered LJP and highlight promising research directions for LFP.
@@
1Junkai Liuthanks: These authors contributed equally to this work and share first authorship., 3Yujie Tong11footnotemark: 1, 1Hui Huang, 1Bowen Zheng, 4Yiran Hu, 3Peicheng Wu, 2Chuan Xiao, 2Makoto Onizuka, 1Muyun Yangthanks: These authors share corresponding authorship: Muyun Yang <yangmuyun@hit.edu.cn>, Shuyuan Zheng <zheng@ist.osaka-u.ac.jp>., 2Shuyuan Zheng22footnotemark: 2 1Harbin Institute of Technology, 2The University of Osaka, 3Zhejiang University, 4Tsinghua University
footnotetext: Our code and data are available at https://github.com/teijyogen/LFP.
@@ 1 Introduction
Extensive research efforts have been devoted to achieving accurate LJP. However, existing LJP research is mostly limited to fact-based LJP (luo2017learning; zhong2018legal; chen2019charge; yue2021neurjudge; feng2022legal; wu2022towards; gan2023exploiting), where the input to the LJP system consists of (ground-truth) legal facts, i.e., facts that are formally established through evidence and determined by the judge. However, the users of LJP, such as litigants and lawyers, typically confirm their legal facts at a very late stage of the litigation medvedeva2023legal. Consequently, the application of LJP is largely confined, as the users often seek to predict judgments before litigation or in its early stages to develop and adjust litigation strategies or related plans.
@@
To further facilitate research on LFP and LFP-empowered LJP, this paper introduces the first benchmark dataset for LFP, LFPBench, which contains evidence items, legal facts, and judgment outcomes collected from 657 litigation cases in China, covering 10 representative types of civil cases. As such, it can be used to evaluate both the LFP and LJP tasks.
@@
  • First, we propose a novel task, legal fact prediction (LFP), which empowers LJP applications to operate in a wider range of real-world scenarios.
@@
Table 1: Comparison between LFPBench and existing LJP benchmarks.
Benchmark
Evidence
items
Claims
Legal
facts
Textual
judgments
Judgment
labels
Label
classes
Type of cases Jurisdiction
SwissJP niklaus2021swiss 2 Generic Switzerland
LJP-MSJudge ma2021legal 3 Civil Mainland China
CAIL2018 xiao2018cail2018 2 Criminal Mainland China
ILDC malik2021ildc 2 Generic India
Auto-Judge long2019automatic 2 Civil Mainland China
BrCase bertalan2020predicting 2 Generic Brazil
PhilCases virtucio2018predicting 2 Criminal Philippines
LFPBench (ours) 3 Civil Mainland China
@@ 2.1 Background
In the legal context, evidence refers to any material or information used to make the existence of a fact more or less probable LII:Evidence, whereas legal facts, also known as findings of fact, are the facts of a case determined by the judge during litigation, based on the presentation and cross-examination of evidence by the parties in a trial LII:FindingOfFact. In other words, only facts that can be substantiated by evidence in a court of law can be acknowledged by the judge as legal facts.
As depicted in Figure 2, in civil law countries such as Germany, France, and China, as well as in common law countries like the UK and the US, a trial primarily resolves the following two tasks to reach a judgment:
@@
As evident, legal facts serve as the foundation for the application of law, and before legal fact-finding is complete, it is logically impossible to predict a judgment based on legal facts. In fact, legal facts are usually finalized when the judge reaches a judgment. Therefore, as medvedeva2023legal has pointed out, utilizing ground-truth legal facts for LJP is impractical. Instead, litigants typically complete evidence collection before litigation or in its early stages, making evidence-based LJP a more reasonable choice. However, existing research on LJP primarily assume the accessibility of legal facts, namely they are all limited to fact-based LJP (e.g., (luo2017learning; zhong2018legal; chen2019charge; yue2021neurjudge; feng2022legal; wu2022towards)), which is mismatched with real-world legal practice.
@@ 2.2 Task Definition
Note that the LFP task is not to summarize the evidence into legal facts. Instead, the available evidence items represent fragmented pieces of legal facts rather than a complete picture. Therefore, the LFP task involves deducing and expanding the evidence into legal facts. Moreover, conflicts or contradictions may exist among evidence items, particularly between the evidence presented by the plaintiff and the defendant. This requires the LFP system to assess the strength and logical coherence of the evidence to resolve the conflicts, making the prediction of legal facts a significant challenge. Notice the plaintiff’s claims and the evidence list are typically available before the trial; therefore, leveraging this information as input for LFP aligns with real-world legal practice. We discuss the flexible adaptability of the input across different scenarios in Appendix A.
@@
Figure 4: Distribution of case types in the LFPBench dataset.
3 LFPBench
@@ 3.1 Dataset Overview
As shown in Table 1, existing LJP datasets, such as CAIL2018 xiao2018cail2018 and ILDC malik2021ildc, focus on the prediction tasks of prison terms and charges rather than including the evidence items submitted by litigants, making LFP infeasible cui2023survey. Therefore, we propose the first benchmark dataset for the LFP task, LFPBench. LFPBench consists of data for 657 first-instance cases in China, covering ten types of civil causes of action as shown in Figure 4. Each case includes the plaintiff’s claims, the evidence items submitted by the litigants, ground-truth legal facts, ground-truth judgments for the claims, and more. Therefore, LFPBench can be used for evaluating both LFP and LJP tasks. We have selected some widely used datasets for comparison with LFPBench, as shown in the Table 1. Not only does LFPBench include a third category for partial support by the court (the other two works only include support or opposition), but the input length also far exceeds theirs, posing a higher challenge to the model’s capabilities. More importantly, to the best of our knowledge, LFPBench is the only dataset that considers predictions made before the trial in real-world scenarios, thus featuring a unique input of evidence lists, while other benchmark datasets follow the paradigm of using legal facts for judgment prediction.
The data statistics of LFPBench can be found in Table 2. In LFPBench, defendants present counterevidence in 58.9% of cases, leading to disputes over the determination of legal facts. Consequently, only 38.71% of claims are fully supported, and plaintiffs completely win only 25.27% of the cases. Therefore, predicting the legal facts and judgments of these cases is highly challenging. Figure 3 presents a data sample from a house lease case, illustrating how conflicting evidence between the litigants complicates the determination of legal facts.
@@ 3.2 Dataset Construction
LFPBench data was extracted from judicial judgments in China. Our legal experts selected ten representative civil litigation causes of action with a moderate level of difficulty in establishing legal facts and retrieved 100 written judgments for each type from the China Judgments Online database Wenshu:CourtDocumentsChina. These case types encompass various common disputes over property and personal rights in daily life. To ensure quality, three legal experts reviewed the judgments and excluded those with overly vague descriptions of evidence. Ultimately, 657 cases were retained.
Then, we used regular expressions to extract legal facts and judgment outcomes, as they are typically written in a consistent structure in the documents. Conversely, since the writing format for claims and evidence varies across judgments, we employed GPT-4o to extract this information. Afterward, our legal experts conducted a manual review to ensure consistency among the extracted claims, evidence, legal facts, and judgments.
Finally, our legal experts annotated the judgment outcomes. Specifically, as shown in Figure 3, they aligned each claim with its corresponding judgment and categorized the outcome into three labels based on the level of support: fully supported, partially supported, and rejected. These labels enable us to evaluate the LFP task using classification-based assessment methods.
Table 3: Human evaluation of the quality of the extracted evidence in LFPBench, conducted by two legal experts. For comparison, the combined rates of affirmation and withdrawal in second-instance civil cases in China across different years are presented.
Metric Value
Accuracy of Evidence-Based LJP
by Legal Experts
87.62%
Affirmation + Withdrawl Rate in 2022 74.78%
Affirmation + Withdrawl Rate in 2023 75.61%
Affirmation + Withdrawl Rate in 2024 76.63%
3.3 Human Evaluation for Extracted Evidence
In many major jurisdictions such as China, Germany, France, and Japan, access to original evidentiary materials is, in principle, restricted to the parties involved in the case due to privacy considerations. Therefore, we opted to extract evidentiary information from publicly available judicial documents. Although such extracted evidence may lack some details compared to the original materials, the purpose and core content of the evidence are generally faithfully reflected in the judgment texts.
Additionally, we asked our two legal experts to perform evidence-based LJP using our dataset, i.e., to predict the judgment outcomes solely based on the extracted evidence. They were required to evaluate a randomly selected sample of 100 cases, comprising a total of 259 plaintiff claims. As shown in Table 3, the experts achieved an LJP accuracy of 87.26%. This level of accuracy is notably high, considering that real-world judicial decisions are not entirely error-free. According to statistics released by the Supreme Court of China, the combined rates of affirmation and withdrawal in second-instance civil cases nationwide were 74.78%, 75.61%, and 76.63% in 2022, 2023, and 2024, respectively Gongbao:JudicialStatsChina. These findings suggest that our extracted evidence retains the vast majority of critical information, enabling legal experts to make accurate judgments accordingly.
@@ 3.3 Human Evaluation for Extracted Evidence
Table 5: LFP similarities under different models.
ROUGE ChatLaw LLM-as-Judge
GPT-4o 0.1808 0.7629 5.52
Claude3.5 0.2138 0.7668 5.83
Qwen2.5-14B 0.1692 0.7464 5.67
Llama3.1-Chinese-8B 0.1763 0.7549 6.23
LawLlama3.1-8B 0.1785 0.7455 5.46
LawJustice-8B 0.1721 0.7069 3.70
@@ 4.1 Setup
  • RQ1 (Model & LJP Approach Comparison): How do SOTA models perform on the LFP and LJP tasks? How do different LJP approaches, including evidence-based LJP, fact-based LJP, and LFP-empowered LJP, perform?
  • RQ2 (Challenge & Bias Analysis): What are the challenges of the LFP task? What biases do existing models exhibit when performing LFP?
@@
We employ 6 LLMs as the LFP and LJP systems, including the closed-source, general-purpose LLMs GPT-4o gpt4o and Claude-3.5-Sonnet-20241022 Anthropic:ClaudeModels, the open-source, general-purpose LLMs Qwen2.5-14B-Instruct yang2024qwen2 and Llama3.1-Chinese-8B llama3.1, as well as the open-source legal LLMs LawJustice-Llama3.1-8B lawjustice and Law-Llama3.1-8B llamalaw. Additionally, we evaluated other legal LLMs including DISC-LawLLM yue2024lawllm, Lawyer-Llama-13B-V2 Lawyer-LLama and AIE-51-8-Law-Model AIE. However, these models failed to perform the LFP task due to poor instruction-following capabilities, as detailed in Appendix C.2.
@@
We compare the following LJP approaches that differ in their input, with their prompts detailed in Appendix C.4.
@@
Figure 5: The correlation between the LFP similarity and the LJP accuaracy. We leverage the DP-Prompt method utpala2023locally to generate rewritten legal facts with varying LFP similarities.
@@
Metrics
We primarily use LJP accuracy as the metric to quantify the influence of LFP on LJP. To measure LFP similarity—that is, the similarity between predicted and ground-truth legal facts—we employ three metrics: (1) ROUGE lin2004rouge; (2) ChatLaw-based similarity, which calculates the distance between the predicted and ground-truth facts based on their embeddings determined by the ChatLaw-Text2Vecw model cui2023chatlaw; and (3) LLM-Judge scores assigned by GPT-4o on a ten-point scale. We also report F1 scores in Appendix C.3, which demonstrate consistency across metrics.
Table 8: The proportions of cases in which each model predicts a complete win/partial win/loss for the plaintiff using LFP-empowered LJP. Evidence items are ordered, ensuring all evidence from the plaintiff (or defendant) appears first.
@@
4.2 Model & LJP Approach Comparison (RQ1)
@@
As shown in Table 4, the closed-source models GPT-4o and Claude3.5 consistently achieve the best performance across different LJP approaches, while the open-source legal LLMs Law-Llama3.1-8B and LawJustice-Llama3.1-8B perform the worst, with accuracy close to random guessing. For the LFP performance in Table 5, the open-source legal LLMs again perform the worst, with the lowest performance among all metrics. This discrepancy cannot be attributed solely to the relatively small size of the legal LLMs, as Llama3.1-Chinese-8B, which has the same model size, performs significantly better. One possible explanation is that these legal LLMs are typically fine-tuned on short-text legal QA datasets lawsft1; yue2024lawllm; Lawyer-LLama, making them less capable of handling complex tasks such as LFP and LJP, which require summarization, reasoning, and deduction over long texts. Future research could explore incorporating general-domain instruction data into fine-tuning to mitigate the catastrophic forgetting of fundamental capabilities. It is also promising to develop more complex reasoning datasets in the legal domain.
@@
As shown in Table 4, for the general-purpose LLMs, predicting legal judgments directly from evidence results in a $\frac{50.46-46.72}{50.46}=7.42\%$ decrease in accuracy compared to predictions based on ground-truth legal facts. This indicates that while fact-based LJP research has made considerable progress, its effectiveness heavily relies on the accessibility of legal facts. On the other hand, although the more practice-aligned evidence-based LJP underperforms [^1] Note that for the legal LLMs, evidence-based LJP outperforms the other approaches. However, given their poor performance close to random guessing, this difference is likely due to randomness rather than a meaningful advantage., the incorporation of LFP reduces the performance gap by 38.50%. Therefore, predicting legal facts from evidence first and then making legal judgments based on the predicted facts can significantly improve the accuracy of evidence-based LJP. Compared to fact-based LJP, LFP-empowered LJP accommodates a broader range of LJP scenarios in legal practice, striking a favorable balance between accuracy and applicability.
@@
Using the DP-Prompt method utpala2023locally, we rewrite the predicted legal facts with the Qwen2.5-14B-Instruct model, generating four versions with varying levels of LFP similarity. We then perform LJP on each version of the rewritten facts and repeated the experiment three times for each parameter setting. As shown in Figure 5, there is a positive correlation between the similarity of the rewritten legal facts and the accuracy of the corresponding legal judgments. These results further underscore the importance of the LFP task in enhancing LJP performance: more accurate legal facts lead to more accurate legal judgments.
4.3 Challenge & Bias Analysis (RQ2)